# FastVID: Dynamic Density Pruning for Fast Video Large Language Models

### v1

@@ 4.3 Ablation Study
Ablation study on $c$ and $\tau$ in DySeg. We evaluate the effect of $c$ and $\tau$ in Eq. (2) of DySeg. Figure 5(a) shows results with varing $c$. In simple video scenarios with high transition similarity, $c$ regulates segmentation. When $c=1$, segmentation relies solely on $\tau$. $c=32$ divides the video into single-frame segments. Performance remains stable for $1\leq c\leq 16$, but the decline at $c=32$ suggests that excessive segmentation disregards temporal relationships. The optimal performance at $c=8$ indicates the benefit of a minimum cluster number. Figure 5(b) shows results with varing $\tau$. When $\tau$ is small, most transitions are selected by $c$. As $\tau$ increases, more transitions fall below the threshold. The best performance is achieved at $\tau=0.9$, effectively grouping redundant frames while separating non-redundant ones.
@@
Ablation study on $p$ and $\beta$ in DTM. In Figure 6(a), we evaluate the effect of $p$ on anchor frame selection. Anchor tokens are sampled from anchor frames. When $p=1$, anchor tokens are evenly distributed across all frames within a segment. As $p$ increases, tokens concentrate on fewer frames. Given the high similarity between frames in a segment, a sufficient number of anchor tokens per anchor frame is essential to capture spatial context. However, if $p$ is too large, all anchor tokens come from the first frame, limiting temporal information. We find that $p=4$ yields the best performance by effectively capturing spatiotemporal context. Figure 6(b) shows the effect of $\beta$ in Eq. (5). When $\beta=0.0$, matching tokens are averaged. $\beta=1.0$ discards all non-anchor tokens, removing segment visual context and leading to a performance drop. Notably, Anchor-centric Assignment (when $\beta=0.6$) yields optimal results, highlighting the importance of representative tokens and visual context.
@@
5 Conclusion
In this paper, we introduced FastVID, a novel inference-time pruning framework designed to accelerate Video LLMs by effectively reducing spatiotemporal redundancy. Through a comprehensive analysis of video tokens from both temporal context and visual context, FastVID dynamically partitions videos into temporally ordered segments and employs density-based token pruning within each segment. This approach significantly decreases computational overhead while maintaining crucial temporal and visual integrity. Extensive experiments across different Video LLMs and diverse video benchmarks confirm the effectiveness and strong generalization capability of our method.

### v2

@@ B.1 Ablation study on $c$ and $\tau$ in DySeg
We evaluate the effect of $c$ and $\tau$ in Eq. (2) of DySeg. Figure 6(a) shows results with varing $c$. In simple video scenarios with high transition similarity, $c$ regulates segmentation. When $c=1$, segmentation relies solely on $\tau$. $c=32$ divides the video into single-frame segments. Performance remains stable for $1\leq c\leq 16$, but the decline at $c=32$ suggests that excessive segmentation disregards temporal relationships. The optimal performance at $c=8$ indicates the benefit of a minimum cluster number. Figure 6(b) shows results with varing $\tau$. When $\tau$ is small, most transitions in $\mathbf{S}$ are selected by $c$. As $\tau$ increases, more transitions fall below the threshold. The best performance is achieved at $\tau=0.9$, effectively grouping redundant frames while separating non-redundant ones.
B.2 Ablation study on $p$ and $\beta$ in DTM
@@
In Figure 7(a), we evaluate the effect of $p$ on anchor frame selection, from which anchor tokens are subsequently sampled. When $p=1$, anchor tokens are evenly distributed across all frames within a segment. As $p$ increases, fewer frames are selected as anchors, while more anchor tokens are allocated to each anchor frame. Given the high similarity between frames in a segment, a sufficient number of anchor tokens per anchor frame is essential to capture visual context. However, if $p$ is too large, all anchor tokens come from the first frame, limiting temporal information. We find that $p=4$ yields the best performance by effectively capturing spatiotemporal context. Figure 7(b) shows the effect of $\beta$ in Eq. (5). When $\beta=0.0$, matching tokens are averaged. $\beta=1.0$ discards all non-anchor tokens, removing segment visual context and leading to a performance drop. Notably, Anchor-centric Assignment (when $\beta=0.6$) yields optimal results, highlighting the importance of representative tokens and visual context.
@@
B.3 Additional Qualitative Results
Figures 8-10 show additional qualitative results on VideoMME. We use the following settings: $c=0,\tau=0.9,d=0.4,p=2$, retaining an average of 10 tokens per frame. Figure 8 presents static scenes, while Figures 9 and 10 present dynamic scenes. These visualizations demonstrate FastVID’s ability to adaptively segment videos across varying temporal dynamics. Within each segment, the merged region boundaries loosely align with object shapes, highlighting the effectiveness of the proposed DTM in preserving visual context.
Table 7: Comparison of state-of-the-art methods on Qwen2-VL [37]. A/B in # Tokens indicates that A tokens are provided as input to the LLM, and subsequently compressed to B tokens during the LLM forward pass. FastV performs pruning based on LLM attention scores, but its eager-attention implementation materializes the full attention matrix in memory, leading to OOM errors due to the large number of video tokens.
Method # Token TFLOPs VideoMME
Short Medium Long Overall
Vanilla 13447.1 100% 124.04 100% 74.1 60.4 54.3 63.0 100%
FastV [5]${}_{\text{ECCV'24}}$ 13447.1/3361.8 100%/25% 31.34 25.3% Out of Memory
VisionZip [42]${}_{\text{CVPR'25}}$ 3349.3 24.9% 24.11 19.4% 70.9 56.3 48.3 58.5 92.9%
PruneVID [13]${}_{\text{ACL'25}}$ 3460.6 25.7% 24.99 20.1% 66.7 54.0 48.1 56.3 89.4%
FastVID 3349.3 24.9% 24.11 19.4% 72.7 58.3 50.7 60.6 96.2%
B.4 Results on Qwen2-VL
Unlike LLaVA-OneVision [16] and LLaVA-Video [47], Qwen2-VL [37] employs a distinct architecture that samples 768 frames per video and processes them using 3D convolutions. It introduces M-RoPE, which decomposes rotary embeddings into temporal, height, and width components. The model dynamically adjusts the resolution of each frame. In Table 7, our FastVID outperforms other baselines on Qwen2-VL under similar model complexity. In particular, FastVID achieves a 2.4 accuracy improvement on the long subset. As shown in Tables 1, 2, and 7, our FastVID’s consistent superiority across three distinct Video LLM architectures further validates its strong generalizability and practical effectiveness.
Appendix C Additional Discussions
C.1 Limitations
FastVID achieves strong performance on LLaVA-OneVision [16] (32 frames/video), maintaining comparable accuracy with only 8.3% of the FLOPs. However, on LLaVA-Video [47] (64 frames/video) and Qwen2-VL [37] (768 frames/video), although FastVID consistently outperforms existing SoTA baselines, a noticeable accuracy drop occurs compared to the original model. We hypothesize that this is primarily due to the increased number of sampled frames, which leads to fewer tokens per frame and increases the semantic importance of each token. Consequently, aggressive pruning risks discarding critical information. In addition, although our query-agnostic token pruning effectively reduces spatiotemporal redundancy, it fails to emphasize query-relevant information sparsely distributed across extensively sampled frames. To overcome these challenges, future work may explore integrating text-guided keyframe selection to better support long-frame Video LLMs [37, 47].
C.2 Broader Impacts
FastVID accelerates inference for existing Video LLMs without modifying their parameters or training new models, thereby minimizing the risk of introducing new biases or unintended behaviors. However, FastVID inherits any potential negative societal impacts of the original models, such as representational bias or potential misuse. Despite this, FastVID maintains strong performance even under extreme token compression. This makes the use of Video LLMs more practical in environments with limited computational resources, promoting broader and more sustainable deployment.
Figure 8: Segment boundaries are marked by brown vertical lines. Tokens generated by DTM and ATS are highlighted in red and blue, respectively. Anchor frames, indicated by red boxes, are processed by DTM individually. In DTM, patches with matching inner and border colors are merged.
Figure 9: Segment boundaries are marked by brown vertical lines. Tokens generated by DTM and ATS are highlighted in red and blue, respectively. Anchor frames, indicated by red boxes, are processed by DTM individually. In DTM, patches with matching inner and border colors are merged.
Figure 10: Segment boundaries are marked by brown vertical lines. Tokens generated by DTM and ATS are highlighted in red and blue, respectively. Anchor frames, indicated by red boxes, are processed by DTM individually. In DTM, patches with matching inner and border colors are merged.